If a marble is dropped from a bridge and reached the water at 2 seconds.
What is its speed before it reaches the water and the height of the bridge?
1
Expert's answer
2018-09-17T15:32:08-0400
The speed of a marble before it reaches the water
v_f=v_i+gt=0+9.8×2=19.6 m/s
The height of the bridge
h=v_i t+(gt^2)/2=0×2+(9.8×2^2)/2=19.6 m